# Discrete-Time CRLB-based Power Allocation for CF MIMO-ISAC with Joint Localization and Velocity Sensing

### v2

@@ VII-B Power Allocation for ISAC
The performance of the proposed PA algorithms is evaluated by comparison with the PP-NCG and PP-NSD benchmark algorithms. To ensure communication quality, a larger number of APs are designated for transmitting ISAC signals, while a smaller subset is used for receiving target-reflected signals. Without loss of generality, we consider a CF MIMO-ISAC system comprising 10 transmitters and 2 receivers, as illustrated in Fig. 2.
To guarantee the CRLB effectiveness, we assume the minimum PA coefficient vector is ${\bm{\rho}}_{\min}=0.01\times{\bm{1}}_{N}$. The PA budget is ${\bm{\rho}}_{\max}=0.3\times{\bm{1}}_{N}$. The SENR is set to be $-10$ dB. Within the coherence time for RCS, a typical modulus-squared RCS matrix is given by $\begin{bmatrix}0.37&0.70&1.38&0.65&2.40&0.18&0.25&0.82&0.14&0.35\\ 0.40\!&\!2.05&1.14&0.42&0.03&0.11&2.02&1.65&1.24&3.65\end{bmatrix}^{\rm T}$. Similarly, the channel gains between the user and the transmitters are assumed to be ${\bm{g}}=[2.11,12.57,5.63,0.75,0.61,1.75,0.20,2.34,14.79,9.68]^{\rm T}$. We consider the OCDM waveform with $M=16$, $T=10^{-2}$ s and a sampling rate of $f_{s}=10$ kHz. Under the given ISAC system setup and waveform configuration, the CRLB thresholds for position and velocity are set to $[250,0.13]$.
@@
The PA vector is initialized as the uniform PA vector ${\bm{\rho}_{1}}={\bm{1}}_{N}/N$, unless noted otherwise. The initial penalty factor is set to be $\mu_{1}=10^{4}$. We hereafter consider the ILS controlling factor $\epsilon_{L}=1e-3$, the initial step size controlling factor $\epsilon_{\upsilon}=0.9$, the penalty factor scaling coefficient $\phi=10$, and the penalty threshold $\epsilon_{\mu}=10^{-3}$, unless noted otherwise. The other simulation conditions are given in Table I.
We begin by evaluating the convergence and steady-state performance of the proposed algorithms. The SINR results corresponding to the iterative PA vectors for all four algorithms are depicted in Fig. 4. As observed, the PP-MCG-ILS algorithm exhibits the fastest convergence, followed by the PP-MSD-ILS. In contrast, the benchmark algorithms converge significantly more slowly. Furthermore, the benchmarks without ILS display noticeable numerical fluctuations, particularly under large fixed step sizes, suggesting potential instability and the risk of divergence. To ensure convergence reliability, we refrain from further increasing the step sizes in these simulations.
[…]
@@
Figure 5: PA coefficient: (a) PP-MCG-ILS; (b) PP-NCG with $\upsilon=2\times 10^{-5}$.
Figure 6: CRLB values: (a) PP-MCG-ILS; (b) PP-NCG with $\upsilon=2\times 10^{-5}$.
The sequential penalty-based algorithms above leverage the optimal solution from the previous iteration loop as the initial point for the current loop under an updated penalty factor. To further explore the impact of initialization, we consider fixing the penalty factor and comparing the ISAC performance under different PA initialization strategies. In addition to the commonly used uniform PA initialization, we introduce a heuristic PA scheme, where the initial allocation is proportional to the communication channel gain, given by ${\bm{\rho}}_{1}={\bm{g}}/\sum_{n}|g_{n}|^{2}$. The results are illustrated in Figs. 7 and 8. When comparing to sequential penalty-based counterparts in Fig. 4, the PP-MCG-ILS and PP-NCG algorithms initialized with uniform PA exhibit similar convergence behavior. However, the PP-MSD-ILS and PP-NSD algorithms show degraded convergence, highlighting the superior convergence properties of CG-based algorithms over SD-based methods under identical conditions. When employing the heuristic PA initialization, all algorithms demonstrate significantly better convergence rates compared to the uniform PA case. Moreover, both PP-NCG and PP-NSD achieve noticeable performance gains relative to their sequential penalty-based counterparts, particularly PP-NCG. In contrast, the PP-MCG-ILS algorithm shows robust and superior convergence behavior, and PP-MSD-ILS exhibits a slight degradation, relative to their sequential penalty-based counterparts.
@@
Figure 9: Steady-state SINRs and CRLBs versus the location CRLB thresholds: (a) SINR; (b) CRLB.
Figure 10: Steady-state SINRs and CRLBs versus the velocity CRLB thresholds: (a) SINR; (b) CRLB.
To summarize, the proposed algorithms exhibit superior convergence behavior under the sequential penalty framework, consistently outperforming the benchmark methods. While the benchmark algorithms can benefit from specific conditions, such as an empirically tuned penalty factor and a well-initialized starting point, their performance is less robust and generalizable. In contrast, the PP-MCG-ILS algorithm demonstrates consistently strong convergence and steady-state accuracy across a wide range of settings, irrespective of initialization or penalty factor selection.
We now investigate the impact of the CRLB thresholds on the performance of the proposed PP-MCG-ILS algorithm. As shown in Fig. 9, relaxing the location CRLB thresholds leads to a gradual increase in the steady-state SINR, which eventually stabilizes. This trend aligns with the behavior observed in Fig. 9, where the achieved location CRLB value closely approaches its threshold and then converge. These trends reveal the trade-off between the communication SINR and the location CRLBs. Furthermore, it is noteworthy that the velocity CRLB remains largely unaffected by variations in the location CRLB thresholds. In contrast, Fig. 10 shows that the performance of the PP-MCG-ILS algorithm exhibits negligible variation with the increased velocity CRLB threshold. The observed CRLB saturation with respect to the corresponding CRLB threshold may be an inherent outcome of the introduced power budget constraints.
@@ VII-C Power Allocation for Pure Sensing
We now verify the effectiveness of the proposed P-MCG-ILS algorithm for total power minimization in pure sensing scenario with joint location and velocity estimation. The initial penalty factor is set to be $\mu_{1}=1$, $\sigma_{\upsilon}=0.2$ and the penalty threshold $\epsilon_{\mu}=10^{-6}$. The other simulation conditions are the same as the PP-MCG-ILS algorithm in Fig. 4.
The iterative results for total power optimization and the corresponding CRLB performance are illustrated in Fig. 11. As seen in Fig. 11, the algorithm demonstrates rapid convergence. Under the same CRLB threshold settings, the total power consumption in the pure sensing scenario is significantly lower than that in the ISAC case. This is expected, as ISAC systems must allocate sufficient power to meet both sensing and communication requirements, particularly for maximizing communication performance. Moreover, both the location and velocity CRLBs in the pure sensing case closely align with their respective thresholds. In contrast, in the ISAC scenario (refer to Fig. 6), the velocity CRLB matches its threshold loosely. This difference arises because, in pure sensing, any enhancement in CRLB performance directly benefits from increased sensing power. However, in ISAC, PA is optimized primarily for communication performance, with CRLB constraints being satisfied, potentially loosely, within acceptable margins.
@@ VIII Conclusions
This paper presents, for the first time, a comprehensive investigation of the power allocation problem in a CF MIMO-ISAC framework, where joint location and velocity estimation were considered as the sensing task. We employed newly derived approximate CRLBs as sensing performance metrics, alongside user receive SINR as the communication metric. To address the inherent nonlinearity and non-convexity of the resulting optimization problem, we propose the PP-MCG-ILS algorithm. Additionally, the P-NCG-ILS algorithm was further proposed to handle the power minimization problem in pure sensing scenarios. We devoted into the detailed analysis of the convergence behavior and qualitatively compare the computational complexity of all proposed methods. Simulation results confirmed the accuracy of the derived CRLBs and demonstrate the efficacy of the proposed power allocation strategies in enhancing pure sensing and overall ISAC performance. Our findings offer valuable insights into resource-efficient design for next-generation CF MIMO-ISAC networks and lay the groundwork for future extensions involving more complex sensing tasks and dynamic resource management.
@@ Appendix A Relevant parameters of the derived CRLBs
where $S$ denotes the number of samples. In addition, the remaining parameters, including $\beta_{n,k}$, $\zeta_{n,k}$, $\xi_{n,k}$, $\varrho_{n,k}$, $\eta_{n,k}$, and $\kappa_{n,k}$, represent geometric spreading factors, given respectively by
$[\beta_{n,k},\zeta_{n,k}]^{\rm T}=$ $\dfrac{|\alpha_{n,k}|}{c}\left(\dfrac{{\bm{l}}-{\bm{l}}_{n}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}+\dfrac{{\bm{l}}-{\bm{l}}_{k}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\right),$ (62a)
$[\xi_{n,k},\varrho_{n,k}]^{\rm T}=$ $\dfrac{f_{c}|\alpha_{n,k}|}{c}\left(\dfrac{{\bm{l}}_{n}-{\bm{l}}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}+\dfrac{{\bm{l}}_{k}-{\bm{l}}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\right),$ (62b)
$[\eta_{n,k},\kappa_{n,k}]^{\rm T}=$ $\dfrac{f_{c}|\alpha_{n,k}|}{c}\bigg{(}({\bm{l}}_{n}-{\bm{l}})\dfrac{{\bm{v}}^{\rm T}({\bm{l}}_{n}-{\bm{l}})}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}^{3}}-\dfrac{{\bm{v}}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}$
@@ Appendix C Proof of Theorem 1
In light of (42), (44) and (63), the ILS condition requires
@@
where positive ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}$ is considered. Assuming the Lipschitz gradient continuity of $L({\bm{\rho}})$, $\|\nabla^{2}L({\bm{\rho}}_{i})\|$ is upper bounded. This indicates a positive and finite ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}$ and thus non-zero $\upsilon_{i}$. Furthermore, in light of (44), we have
@@
Since $L({\bm{\rho}})$ is a monotonously decreasing and bounded function regarding ${\bm{\rho}}_{i}$, the (conjugate) gradient-based descent algorithm converges to $\Delta^{2}L({\bm{\rho}}_{i+1})\rightarrow 0$ when $i>I$ for a large constant $I$. However, the deflection factor computed by (41) or the descent direction vector in (49) can guarantee a strict descent ${\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})<0$ at each iteration $i$ unless ${\bm{d}}_{i}={\bm{0}}$. Consequently, sequence $\{L({\bm{\rho}}_{i})\}$ will converge to a stationary point with a zero descent direction ${\bm{d}}_{i}\rightarrow{\bm{0}}$ when it has Lipschitz gradient continuity. According to Lemma 1, we have $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ when ${\bm{d}}_{i}\rightarrow{\bm{0}}$. On the other hand, we assume the objective function is not Lipschitz gradient continuous. In this case, if the sequence $\{{\bm{\rho}}_{i}\}$ pass a Lipschitz noncontinuous point ${\bm{\rho}}_{I}$ where ${\bm{d}}_{I}^{\rm T}\nabla^{2}L({\bm{\rho}}_{I}){\bm{d}}_{I}$ is extremely large, to satisfy (C), the step size should meet $\upsilon_{I}\rightarrow 0$. According to Lemma 1, we have $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ and $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$ when $\upsilon_{i}\rightarrow 0$ for $i\geq I$, i.e., the sequences $\{{\bm{\rho}}_{i}\}$ and $\{L({\bm{\rho}}_{i})\}$ converge to a Lipschitz gradient noncontinuous point. $\blacksquare$

### v3

@@ VII-B Power Allocation for ISAC
The performance of the proposed PA algorithms is evaluated by comparison with the PP-NCG and PP-NSD benchmark algorithms. To ensure communication quality, a larger number of APs are designated for transmitting ISAC signals, while a smaller subset is used for receiving target-reflected signals. Without loss of generality, we consider a CF MIMO-ISAC system comprising 10 transmitters and 2 receivers in Fig. 2.
To guarantee the CRLB effectiveness, we assume the minimum PA coefficient vector is ${\bm{\rho}}_{\min}=0.01\times{\bm{1}}_{N}$. The PA budget is ${\bm{\rho}}_{\max}=0.3\times{\bm{1}}_{N}$. The SENR is set to be 10-10 dB. Within the coherence time for RCS, a typical modulus-squared RCS matrix is given by $\begin{bmatrix}0.37&0.70&1.38&0.65&2.40&0.18&0.25&0.82&0.14&0.35\\ 0.40\!&\!2.05&1.14&0.42&0.03&0.11&2.02&1.65&1.24&3.65\end{bmatrix}^{\rm T}$. Similarly, the channel gains between the user and the transmitters are assumed to be ${\bm{g}}=[2.11,12.57,5.63,0.75,0.61,1.75,0.20,2.34,14.79,9.68]^{\rm T}$. We consider the OCDM waveform with $M=16$, $T=10^{-2}$ s and a sampling rate of $f_{s}=10$ kHz. Under the given ISAC system setup and waveform configuration, the CRLB thresholds for position and velocity are set to $[250,0.13]$.
@@
The PA vector is initialized as the uniform PA vector ${\bm{\rho}_{0}}={\bm{1}}_{N}/N$, unless noted otherwise. The initial penalty factor is set to be $\mu_{1}=10^{4}$. We hereafter consider the ILS controlling factor $\epsilon_{L}=1e-3$, the initial step size controlling factor $\epsilon_{\upsilon}=0.9$, the penalty factor scaling coefficient $\phi=10$, and the penalty threshold $\epsilon_{\mu}=10^{-3}$, unless noted otherwise. The other simulation conditions are given in Table I.
@@
Figure 5: PA: (a) $P$-MCG-ILS; (b) $P$-NCG with $\upsilon=2\times 10^{-5}$.
Figure 6: CRLBs: (a) $P$-MCG-ILS; (b) $P$-NCG with $\upsilon=2\times 10^{-5}$.
We begin by evaluating the convergence and steady-state performance of the proposed algorithms. The SINR results corresponding to the iterative PA vectors for all four algorithms are depicted in Fig. 4. As observed, the PP-MCG-ILS algorithm exhibits the fastest convergence, followed by the PP-MSD-ILS. In contrast, the benchmark algorithms converge significantly more slowly. Furthermore, the benchmarks without ILS display noticeable numerical fluctuations, particularly under large fixed step sizes, suggesting potential instability and the risk of divergence. Therefore, we refrain from further increasing the step sizes in these simulations.
The evolutions of the PA coefficients and the corresponding CRLB values over iterations are illustrated in Fig. 5 and Fig. 6, respectively. Overall, the results demonstrate that the PP-MCG-ILS algorithm achieves significantly faster convergence than the PP-NCG algorithm. As shown in Fig. 5, all algorithms eventually converge to the same PA vector, given by ${\bm{\rho}}=[0.01,\ 0.30,\ 0.0638,\ 0.01,\ \dots,\ 0.01,\ 0.30,\ 0.2762]^{\rm T}$, which satisfies the power constraints. Notably, transmitters 2, 9, and 11 are allocated more power due to their superior communication channel gains. Furthermore, Fig. 6 dipicts the exploration process of the CRLB values under threshold-based constraints, in particular for the location CRLB, thereby validating the effectiveness of the proposed optimization framework.
The sequential penalty-based algorithms use the optimal solution from the previous loop as the initialization for the next loop with an updated penalty factor. To further examine the impact of initialization, we fix the penalty factor and compare the ISAC performance under different PA initialization strategies. In addition to uniform initialization, a heuristic scheme is considered, where the initial PA is set proportional to the communication channel gain, i.e., ${\bm{\rho}}_{0}={\bm{g}}/\sum_{n}|g_{n}|^{2}$. The results in Figs. 7 and 8 show that a smaller penalty factor ($\mu=10^{5}$) yields faster convergence than $\mu=10^{6}$, especially for the benchmark algorithms. Moreover, heuristic PA initialization significantly accelerates convergence for all methods compared to uniform initialization. When comparing the fixed-penalty schemes with uniform initialization (Fig. 7) to the sequential penalty-based schemes (Fig. 4), the PP-MCG-ILS and PP-NCG algorithms exhibit similar convergence behavior, whereas the PP-MSD-ILS and PP-NSD algorithms converge more slowly. With heuristic initialization (Fig. 8), PP-NCG and PP-NSD achieve noticeable performance gains, while PP-MCG-ILS maintains robust and superior convergence and PP-MSD-ILS shows a slight degradation.
@@
To summarize, while the benchmark algorithms can benefit from specific conditions, such as an empirically tuned penalty factor and a well-initialized starting point, their performance is less robust and generalizable. In contrast, the PP-MCG-ILS algorithm exhibits superior convergence behavior, consistently outperforming the benchmark methods, irrespective of initialization or penalty factor selection.
We now investigate the impact of the CRLB thresholds on the performance of the proposed PP-MCG-ILS algorithm. As shown in Fig. 9, relaxing the location CRLB threshold results in a gradual increase in the steady-state SINR, which eventually saturates. This behavior is consistent with the trend observed in Fig. 9, where the achieved location CRLB closely approaches its prescribed threshold and then converges. These results illustrate the inherent trade-off between communication SINR and location CRLB.
Figure 9: Steady-state SINRs and CRLBs: (a) SINR; (b) CRLB.
Figure 10: Steady-state SINRs and CRLBs: (a) SINR; (b) CRLB.
However, the velocity CRLB in Fig. 9 is largely insensitive to variations in the location CRLB threshold. Similarly, Fig. 10 shows that changing the velocity CRLB threshold has a negligible impact on the performance of the PP-MCG-ILS algorithm. The observed saturation and difference of the CRLB with respect to its threshold is likely a consequence of the nonlinear structure of the CRLB formulation and the compact feasible region imposed by the power constraints.
@@ VII-C Power Allocation for Pure Sensing
We now evaluate the proposed P-MCG-ILS algorithm on total power minimization in pure sensing scenario. The initial penalty factor is set to be $\mu_{1}=1$, $\sigma_{\upsilon}=0.2$ and the penalty threshold $\epsilon_{\mu}=10^{-6}$. The other simulation conditions are the same as the PP-MCG-ILS algorithm in Fig. 4.
The iterative results for total power optimization and the corresponding CRLB performance are illustrated in Fig. 11. As seen in Fig. 11, the algorithm demonstrates rapid convergence. Under the same CRLB threshold settings, the total power consumption in the pure sensing scenario is significantly lower than that in the ISAC case. This is expected, as ISAC systems must allocate sufficient power to meet both sensing and communication requirements, particularly for maximizing communication performance.
Moreover, both the location and velocity CRLBs in the pure sensing case closely align with their respective thresholds. In contrast, in the ISAC scenario (refer to Fig. 6 and Fig. 10), the velocity CRLB matches its threshold loosely. This difference arises because, in pure sensing, any enhancement in CRLB performance directly benefits from increased sensing power. However, in ISAC, PA is optimized primarily for communication performance, with CRLB constraints being satisfied, potentially loosely, within acceptable margins.
@@ VIII Conclusions
This paper presented a comprehensive investigation of the PA problem in a CF MIMO-ISAC framework, where joint location and velocity estimation were considered as sensing tasks. Newly derived approximate CRLBs were employed as sensing performance metrics, together with the user receive SINR as the communication metric. To address the inherent nonlinearity and nonconvexity of the resulting optimization problem, the PP-MCG-ILS algorithm was proposed. In addition, the P-NCG-ILS algorithm was developed to handle the power minimization problem in pure sensing scenarios. We further conducted a detailed convergence analysis and provided a qualitative comparison of the computational complexity of the proposed methods. Simulation results confirmed the accuracy of the derived CRLBs and demonstrated the effectiveness of the proposed PA strategies in enhancing both pure sensing and overall ISAC performance. These findings provide valuable insights into resource-efficient design for next-generation CF MIMO-ISAC networks and lay the groundwork for future extensions involving more complex sensing tasks and dynamic resource management.
@@ Appendix A Relevant parameters of the derived CRLBs
where $S$ denotes the number of samples. The remaining parameters, including $\beta_{n,k}$, $\zeta_{n,k}$, $\xi_{n,k}$, $\varrho_{n,k}$, $\eta_{n,k}$, and $\kappa_{n,k}$, represent geometric spreading factors, given respectively by
$[\beta_{n,k},\zeta_{n,k}]^{\rm T}=\$$\dfrac{|\alpha_{n,k}|}{c}\left(\dfrac{{\bm{l}}-{\bm{l}}_{n}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}+\dfrac{{\bm{l}}-{\bm{l}}_{k}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\right),$ (62a)
$[\xi_{n,k},\varrho_{n,k}]^{\rm T}=\$$\dfrac{f_{c}|\alpha_{n,k}|}{c}\left(\dfrac{{\bm{l}}_{n}-{\bm{l}}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}+\dfrac{{\bm{l}}_{k}-{\bm{l}}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\right),$ (62b)
$[\eta_{n,k},\kappa_{n,k}]^{\rm T}=\$$\dfrac{f_{c}|\alpha_{n,k}|}{c}\bigg(({\bm{l}}_{n}-{\bm{l}})\dfrac{{\bm{v}}^{\rm T}({\bm{l}}_{n}-{\bm{l}})}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}^{3}}-\dfrac{{\bm{v}}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}$
@@ Appendix C Proof of Theorem 1
From (42), (63), and (44), the ILS condition yields
@@
where positive ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}$ is considered. With the Lipschitz gradient continuity of $L({\bm{\rho}})$, $\|\nabla^{2}L({\bm{\rho}}_{i})\|$ is upper bounded. This indicates a positive and finite ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}$ and thus nonzero and bounded $\upsilon_{i}$. Furthermore, in light of (44), we have
@@
Since $L({\bm{\rho}})$ is monotonously decreasing and bounded, we have $\Delta^{2}L({\bm{\rho}}_{i+1})\rightarrow 0$ when $i>I$ for a large constant $I$. This implies ${\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})\rightarrow 0$. However, a strict descent is satisfied with ${\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})<0$ at each iteration $i$ unless ${\bm{d}}_{i}={\bm{0}}$. Consequently, sequence $\{L({\bm{\rho}}_{i})\}$ will converge to a stationary point with a zero descent direction ${\bm{d}}_{i}\rightarrow{\bm{0}}$ when it has Lipschitz gradient continuity. According to Lemma 1, we further have $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ when ${\bm{d}}_{i}\rightarrow{\bm{0}}$. $\blacksquare$